Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2001-083_2001-089

Mission Operations Days: 2001/083 to 2001/089
March 24, 2001 00:00:00 GMT to March 30, 2001 00:00:00 GMT


**********************************************************
Science Data loss on DOY 2001/087. See Non-MODIS Significant Events for details.
**********************************************************

Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - B Side
ComponentState(s)Comment
BlackbodyA Off; B On (290 K)Nominal
Calibration ElectronicsA Off; B OnNominal
Control ProcessorA Off; B OnNominal
Nadir DoorUnlatched, OpenNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA Off; B OnNominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 Off; #3 & #4 OnNominal
Format ProcessorA Off; B OnNominal
Power Supplies#1 Off; #2 OnNominal
PV VIS FPAA Off; B OnNominal
PV NIR FPAA Off; B OnNominal
PV SMIR FPAA Off; B OnNominal
PV LWIR FPAA Off; B OnNominal
PC LWIR FPAA Off; B OnNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (83 K)Nominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA Off; B OnNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA Off; B OnNominal
Flight SoftwareRev BD + 1 set of patches
  • FR Error Counting and Statistics (2 patches)
Nominal
Inhibit Ids SetNoneNominal
TMONs Enabled68Nominal

This Week’s Completed MODIS Activities:

Saturday, March 24, 2001

None

Sunday, March 25, 2001

None

Monday, March 26, 2001

None

Tuesday, March 27, 2001

None

Wednesday, March 28, 2001

None

Thursday, March 29, 2001

None

Friday, March 30, 2001

None

This Week's Scheduled MODIS Activities Not Completed:

None

Upcoming MODIS Events:

Saturday, March 31, 2001

None

Sunday, April 1, 2001

None

Monday, April 2, 2001

None

Tuesday, April 3, 2001

None

Wednesday, April 4, 2001

None

Thursday, April 5, 2001

None

Friday, April 6, 2001

None

Maneuvers:

There will be no MODIS roll maneuver in April because the roll angle would have been too great.

The next roll maneuver will be scheduled in May.

An increase in Solar Flux has been noted. As a result, the FOT has planned for a Drag Make Up Maneuver at 2001/095-19:03:00 (4/5/2001) and expects to schedule a follow up possibly a week later based on the final result on the 2001/095 maneuver and the current Solar Flux levels.

MODIS Anomalies:

There appears to be a problem processing the science data collected during the SRCA activities on Friday, March 2nd, 2001 (day 061). Specifically, this may be related to the change in sector sizes for the Solar Diffuser and SRCA. If this is the cause, it will affect data from 061/14:06:30 to 061/18:21:56 GMT. This problem is not yet well understood.

General Instrument Comments:

MODIS is in Science Mode on the B-side with the SVD and NAD open, operating with the Formatter Resolution Flight Software Patches (FRFSP).

MODIS Telemetry Trends:

Temperature trends are correlating with the solar beta angle.

Non-MODIS Significant Events:

On Wednesday, March 28, 2001 (DOY 2001/087), there was a loss of MODIS Science Data. The cause of the data loss was a combination of a failure in a new build of the online software, a failure during switch over to the previous version of the online software, and additional time issues. Compounding these software problems was a hardware failure of a matrix switch at the receiving ground station. As a result, once science data playback had begun, insufficient time remained to dump and clear the MODIS SSR Buffer. The Buffer filled up, and was disabled prior to the next available contact. During the following contact, the science data was replayed, and the buffer cleared and re-enabled. The times of note for the event are:

2001/087-19:57:59 MODIS SSR Buffer – full, disabled
2001/087-19:59:48 MODIS FR Rate to Day (part of nominal ops)
2001/087-19:57:59 MODIS SSR Buffer – replayed, cleared, enabled

5568 EDU Blocks (2201341 CADUs) of data were lost (as the timing shows, the majority were Day Rate samples).

Limited Life Item Status:

SRCA 10W Lamp #1: 182.5 of 500 hours
SRCA 10W Lamp #2: 141.1 of 500 hours
SRCA 10W Lamp #3: 150.3 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours

SRCA 1W Lamp #1: 557.2 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours

Solar Diffuser Door: 1567 of 3022 Movements
Nadir Aperture Door: 532 of 1316 Movements
Space View Door: 437 of 1316 Movements